Consider now the following lottery: There are infinitely many tickets, of which 4% are a “win”....

  1. Consider now the following lottery: There are infinitely many tickets, of which 4% are a “win”. Every ticket costs $2, and a winning ticket pays out $120.

    (a) Give a lower bound on the probability that playing the lottery will win you money. (b) Find a price such that your losses can exceed this price with probability at most 25%.
Answer #1

Given: The probability of winning is, P (W)= 0.04 The cost price of a ticket is, CP= $2 Wining Amount is, WA= $120 (a) A contCompute the number of tickets bought as follows- P(L) = 0.25 (0.96)* = 0.25 nlog(0.96)=log (0.25) _log(0.25) n= log(0.96) =